OpenStudy (fluttershyk):

@silveralchemist09 Yvonne put $4000 in a savings account. At the end of 3 years, the account had earned $960 in simple interest. A. how much does she have in her account at he end of 3 years B. at what annual simple interest rate did the account grow? show your work. C. How many more dollars would she have in her account if the interest rate were 1% greater? Show your work.

OpenStudy (anonymous):

I = Prn I = interest P = principal... amount invested r = interest rate as a decimal n = time periods a. Interest + principal = amount in the account b. 960 = 4000 x r x 3 960 = 12000 x r solve for r then multiply by 100 to get a percentage c. find I = 4000 x ( r + 1)/100 x 3

OpenStudy (fluttershyk):

@silveralchemist09 ok im confused can we do one at a time like starting with A

OpenStudy (anonymous):

ok a. Interest + principal = amount in the account

OpenStudy (anonymous):

so the answer is A.

OpenStudy (fluttershyk):

ok instead of words can you use the numbers please like for example 854+6653= amount in interest this is like an essay type question so I have to solve for A B and C

OpenStudy (anonymous):

yes correct

OpenStudy (fluttershyk):

ok A.) would be 4000+960=4960

OpenStudy (anonymous):

correct so its not A.

OpenStudy (fluttershyk):

ok now how do I solve B

OpenStudy (anonymous):

b. 960 = 4000 x r x 3

OpenStudy (fluttershyk):

0.08= r or 8% @silveralchemist09 sorry it took a while to respond I had to leave but is this correct

OpenStudy (fluttershyk):

@silveralchemist09 and for C I think its I=4000*0.09*3 I=$1080

OpenStudy (anonymous):

correct

OpenStudy (fluttershyk):

thank you

OpenStudy (anonymous):

welcome